OpenStudy (anonymous):

When buying mulch for your lawn, you find that it is ordered in square yards. However, you have the area measurements of your lawn in square feet, found by multiplying the length of the yard times the width. Demonstrate with an example by first giving the size of the yard in square feet (length and width), and then convert how much in square yards is needed to cover the area in mulch. (Hint: Research to find how many square feet equals 1 square yard). Because lawns may not come out with exact whole-number measurements, give an example that ended in a fractional value. How would this change yo

OpenStudy (anonymous):

nine square feet will equal on square yard because 3 feet is one yard

OpenStudy (anonymous):

suppose your dimensions are 30 feet by 42 feet. then the area in square feet would be 30×42=1260 square feet. but 30 feet is 10 yards (each yard is 3 feet) and likewise 42 feet is 42÷3=14 yards, so the area in square yards is 10×14=140

OpenStudy (anonymous):

and you can see that 1260÷140=9 so in square yards the number is one ninth of what you get in square feet.